# Finite presentability of outer automorphism groups of large-type Artin groups

### v1

@@ 3 Deformation Spaces for Artin Groups
In this section, given a large-type Artin group $A$, we construct a deformation space $\mathcal{D}$. Trees in this space will correspond to visual splittings of $A_{\Gamma}\cong A$ as amalgamated free products of chunks of $\Gamma$. These chunks lack separating simplices in the following sense.
@@
The second condition will help us when finding point stabilisers in the deformation space we introduce. This part could likely be weakened; indeed, our proof still applies if we only ask that $\mathrm{Out}(A_{\Gamma^{\prime}})$ is finitely generated and virtually abelian. Nonetheless, this weakening doesn’t admit any new classes of Artin groups to be proven chunk rigid, and we expect this remains true in general.
@@
The second case is taken from work in progress by Blufstein, Martin and Vaskou.
@@
Note that the graph of groups we built for $A$ is a tree, by Remark 2.12 the quotient of every $A$-tree in $\mathcal{D}_{\Gamma}$ is a tree, and the simplified theory of deformation spaces we considered in the previous section is sufficient for understanding this space of visual splittings.
@@
Consider the defining graph $\Gamma$ as in Figure 1, for simplicity with all edges labelled 3. It has 4 chunks, coloured red, blue, yellow and green. The purple edge is shared by the red and blue chunks.
@@
If $A$ is connected, large-type Artin group with rigid chunks, then the construction of $\mathcal{D}_{\Gamma}$ did not depend on the choices in the construction of $T_{\Gamma}$, nor the choice of $\Gamma$ and isomorphism $A\cong A_{\Gamma}$.
@@
We use Theorem 2.15. Crucially, notice that $L_{T_{\Gamma}}^{-1}(0)$ is the set of elements fixing a vertex, that is precisely the set of elements conjugated into some chunk parabolic subgroup $A_{\Gamma^{\prime}}$. Immediately we see there is no dependence on the choices made in constructing $T_{\Gamma}$.
Moreover, since $A$ has rigid chunks, if there is another defining graph $\Lambda$ and $A_{\Lambda}\cong A$ (fix this isomorphism), $L_{T_{\Gamma}}^{-1}(0)\subseteq L_{T_{\Lambda}}^{-1}(0)$, and thus by interchanging $\Gamma$ and $\Lambda$, $L_{T_{\Gamma}}^{-1}(0)=L_{T_{\Lambda}}^{-1}(0)$. It follows that $\mathcal{D}_{\Gamma}$ also didn’t depend on the defining graph and isomorphism. We henceforth refer to it simply as $\mathcal{D}$.
Now suppose $(T,\Omega)\in\mathcal{D}$, so by Theorem 2.15 $L_{(T,\Omega)}^{-1}(0)=L_{T_{\Gamma}}^{-1}(0)$. Take $\phi\in\mathrm{Aut}(A)$. Since $A$ is chunk rigid, $\phi$ and $\phi^{-1}$ both send $L_{T_{\Gamma}}^{-1}(0)$ into itself, so $\phi$ and $\phi^{-1}$ restrict to a bijection on $L_{T_{\Gamma}}^{-1}(0)$. In particular,
$L_{(T,\Omega\circ\phi)}^{-1}(0)=\phi^{-1}(L_{(T,\Omega)}^{-1}(0))=\phi^{-1}(L_{T_{\Gamma}}^{-1}(0))=L_{T_{\Gamma}}^{-1}(0)=L_{(T,\Omega)}^{-1}(0).$
So $\phi\cdot(T,\Omega)\in\mathcal{D}$ also. It follows that $\mathrm{Aut}(A)$ acts on $\mathcal{D}$, and since inner automorphisms act trivially, this action factors through an action of $\mathrm{Out}(A)$. ∎
We immediately apply Thoerem 2.21, and restrict our attention to the finite dimensional, $\mathrm{Out}(A)$ invariant subcomplex $\mathcal{F}\subseteq\mathcal{D}$ of surviving $A$-trees.
@@
It is enough to show that in the graph of groups we built for $A_{\Gamma}\curvearrowright T_{\Gamma}$, none of the inclusions of edge groups into vertex groups were equalities. Suppose $e=\{u,v\}$ was such that $G_{u}=G_{e}$, then this induces an inclusion $G_{u}\leq G_{v}$. However this induces an inclusion of the set of standard generators generating $G_{u}$ into those generating $G_{v}$, contradicting maximality in the definition of $G_{u}$ subgraph spanned by generators of $G_{u}$ being a chunk.
@@
4 Finitely Presented Outer Automorphism Groups
Our aim for this section is to show that, given $A$ an abstract connected large-type chunk rigid Artin group, $\mathrm{Out}(A)$ is finitely presented. From here, will will fix $A$ as an abstract connected large-type chunk rigid Artin group. We will write $\mathcal{F}$ for subcomplex of surviving trees in the canonical deformation space $\mathcal{D}$, as discussed in the previous section.
We will prove this in two parts. First we will show that $\mathcal{F}/\mathrm{Out}(A)$ is a finite complex. Secondly, we will prove that the simplex stabilisers of $\mathrm{Out}(A)\curvearrowright\mathcal{F}$ are finitely presented.
@@ 4 Finitely Presented Outer Automorphism Groups
Let $A$ be an abstract connected large-type chunk rigid Artin group. Then $\mathrm{Out}(A)$ is finitely presented.
@@
The group $\mathrm{Out}(A)$ acts on the simplicial complex $\mathcal{F}$. Thus, $\mathrm{Out}(A)$ is the fundamental group of the complex of groups $\mathcal{F}//\mathrm{Out}(A)$. By Proposition 4.11 this complex has finitely many vertices, and since the action $\mathrm{Out}(A)\curvearrowright\mathcal{F}$ is simplicial this means there are finitely many simplices. Moreover, the simplex groups are finitely presented by Lemma 4.15. It follows that the presentation for $\mathrm{Out}(A)$ as the fundamental group of this complex of groups is finite [BH13, III.$\mathcal{C}$, Theorem 3.7]. ∎
@@
Suppose $A_{\Gamma}$ is a connected large-type Artin group, and $\Gamma$ is either triangle-free, or has every label at least 6 (i.e. $A_{\Gamma}$ is XXXL-type). Then $\mathrm{Out}(A_{\Gamma})$ is finitely presentable.
@@ 4.1 Co-finiteness
It is convenient for us to regard $\Gamma$-trees as $A_{\Gamma}$-trees with distinguished fundamental domains, but one could equivalently see them as graph of groups decompositions of $A_{\Gamma}$, with isomorphism being isomorphism of graphs of groups.
Lemma 4.5 justifies our definition of $\Gamma$-trees and isomorphism between them: if two trees in $\mathcal{F}$ can be given the structure of isomorphic $\Gamma$-trees (for the same $\Gamma$, so the definition of $\Gamma$-tree isomorphism makes sense), then they are in the same orbit under the action of $\mathrm{Out}(A)$. We start with a technical lemma, which is essentially the observation that $\Gamma$-tree isomorphisms induce a graph-of-groups isomorphism on the $A_{\Gamma}$-trees, and hence there is an induced equivariant isometry.
@@
Suppose $T_{1},T_{2}$ are isomorphic $\Gamma$-trees. Then they are equivariantly isometric: that is there is an isometry $f:T_{1}\rightarrow T_{2}$ such that for all $g\in A_{\Gamma}$ and $x\in T_{1}$, $f(gx)=gf(x).$
@@
By the definition of $\Gamma$-tree isomorphism, there is an isometry between the fundamental domains of the $\Gamma$-trees, $\hat{f}:K_{1}\rightarrow K_{2}$, which respects the edge and vertex groups. We define $f:T_{1}\rightarrow T_{2}$ extending this isometry by such that for every point $gx\in T_{1}$ where $g\in A_{\Gamma}$ and $x\in K_{1}$ (all points can be written thus, since $K_{1}$ is a fundamental domain), $f(gx):=g\hat{f}(x)$. We need to check that this is a well defined, equivariant, and an isometry.
To see that $f$ is well defined, note that if $gx=hy$ where $x,y\in K_{1}$, then it must be that $x=y$ since $K_{1}$ is a strict fundamental domain. Therefore $h^{-1}g\in G_{x}$. It follows that $h^{-1}g\in G_{\hat{f}(x)}$, since $\hat{f}$ is vertex group preserving, so $f(gx)=g\hat{f}(x)=h\hat{f}(x)=f(hx)$ as required.
It is clear that $f$ sends any edge of $T_{1}$ isometrically to an edge of $T_{2}$. This is because any edge of $T_{1}$ can be written $ge$ where $g\in A_{\Gamma}$ and $e\subseteq K$, and $\hat{f}$ is an isometry of $e$ onto an edge, so $f$ isometrically embeds $ge$ to the edge $g\hat{f}(e)$.
Therefore, to check that $f$ is an isometric embedding, it is sufficient to check that it does not identify edges that share an endpoint. Suppose $e_{1}=\{u_{1},v\}$ and $e_{2}=\{v,u_{2}\}$ are distinct edges of $T_{1}$. If the edges are in distinct orbits, then it is not hard to check, using that fact that $K_{1}$ is a strict fundamental domain, that there is $g\in A_{\Gamma}$ such that $g^{-1}(e_{1}\cup e_{2})\subseteq K_{1}$, and the required result follows from the fact that $\hat{f}$ is an isometric embedding. Otherwise, suppose $h^{-1}e_{1}\subseteq K_{1}$, and so $h^{-1}u_{2}=gh^{-1}u_{1}$ where $g\in G_{h^{-1}v}\setminus G_{h^{-1}u_{1}}$ by assumption that $u_{1}$ and $u_{2}$ are distinct. Note that
$G_{h^{-1}v}\setminus G_{h^{-1}u_{1}}=G_{\hat{f}(h^{-1}v)}\setminus G_{\hat{f}(h^{-1}u_{1})}=G_{h^{-1}f(v)}\setminus G_{h^{-1}f(u_{1})}=h^{-1}(G_{f(v)}\setminus G_{f(u_{1})}),$
where the crucial equaity is the first where we use that $\hat{f}$ respects the vertex groups. In particular,
$hgh^{-1}\in G_{f(v)}\setminus G_{f(u_{1})}$
@@
$f(u_{2})=f(hgh^{-1}u_{1})=hg\hat{f}(h^{-1}u_{1})=hgh^{-1}f(u_{1}),$
and $hgh^{-1}\in G_{f(v)}$, so $hgh^{-1}$ takes $f(e_{1})$ to $f(e_{2})$, but $hgh^{-1}\notin G_{f(u_{2})}$, so these edges are distinct as required.
That $f$ is surjective follows immediately from the fact any point in $T_{2}$ is in a translate of the fundamental domain, so can be written as $g\hat{f}(x)=f(gx)$.
Finally, equivariance of $f$ follows almost immediately from the definition. ∎
@@
Let $(T_{i},\Omega_{i})$ ($i\in\{1,2\}$) be two $A$-trees (given by $\Omega_{i}:A\rightarrow\mathrm{Aut}(T_{i})$) in $\mathcal{F}$. Suppose $\Gamma$ is an abstract defining graph, and there are isomorphisms $\psi_{1},\psi_{2}:A_{\Gamma}\rightarrow A$. Suppose that the $A_{\Gamma}$-trees $(T_{i},\Omega_{i}\circ\psi_{i})$ can be seen as isomorphic $\Gamma$-trees (by choice of fundamental domains $K_{i}$).
@@
We claim that $K=\mathrm{Hull}(\{v_{i}\}_{i\in I})$. To see this, suppose for contradiction there was a nonempty subtree of $K$ outside of this hull, attached to a vertex $v\in\mathrm{Hull}(\{v_{i}\}_{i\in I})$ by an edge $e$. Let $G_{v}=A_{S}$ (we make no assumptions beyond it being standard parabolic, which follows from the definition of a $\Gamma$-tree). Now take $w$ an edge or vertex of this extremal subtree. By definition of a $\Gamma$-tree, this has stabiliser $A_{S^{\prime}}$ where $S^{\prime}$ is either a standard generator, or a pair of standard generators spanning an spherical dihedral parabolic subgroup. Take $s\in S^{\prime}$. This generator, viewed as a vertex in $\Gamma$, occurs in a chunk of $\Gamma$. Therefore it fixes the unique vertex $v_{i}$ in $K$ corresponding to this chunk, and the geodesic from $w$ to $v_{i}$. Given that $w$ is a vertex in an extremal subtree, this geodesic passes through $v$. It follows that $s\in A_{S}$. Since $s$ and $w$ were generic, it follows that the stabiliser of every point in the extremal subtree has stabiliser including into $A_{S}$. This contradicts minimality, completing the proof of the claim.
Now we show that there is a uniform bound on the number of vertices in $K$. Since $K=\mathrm{Hull}(\{v_{i}\}_{i\in I})$, we can write $K$ as a union of $|I|^{2}$ geodesics, in particular $K=\bigcup_{i,j\in I}[v_{i},v_{j}]$. Note that $|I|$ is the number of chunks, and depends only on $\Gamma$.
@@
Now suppose $v$ is internal to $\mu$ with valence exactly 2 in $K$, and the call $e_{1},e_{2}$ the adjacent edges. One case is that $v\in\{v_{i}\}_{i\in I}$, but this happens at most $|I|-2$ times. Thus we focus on the other case, so $G_{v}=A_{S}$, where $S$ is either a single standard generator, or two standard generators spanning an spherical dihedral parabolic subgroup.
@@
For expansion, we may have to change $\Gamma$ to some new $\Gamma^{\prime}$ (with $A_{\Gamma^{\prime}}\cong A_{\Gamma}$). This case is more delicate than the previous one, with the tricky part being picking the right $\Gamma^{\prime}$ and fundamental domain on $T^{\prime}$ (once we have done so, checking there is a $\Gamma^{\prime}$ structure is labourious but not difficult).
@@
Suppose $\varphi:T^{\prime}\rightarrow T$ is a collapse of an edge $e=\{u,v\}$ with $G_{u}\leq G_{v}$. Suppose that $T$ is a $\Gamma$-tree. Suppose that $T,T^{\prime}\in\mathcal{F}$ are both surviving. Then there is $\Gamma^{\prime}$ and an isomorphism $\psi:A_{\Gamma^{\prime}}\rightarrow A_{\Gamma}$, such that $T^{\prime}$, when viewed as an $A_{\Gamma^{\prime}}$-tree (by precomposing the action $A_{\Gamma}\rightarrow\mathrm{Aut}(T^{\prime})$ with $\psi$) is a $\Gamma^{\prime}$-tree.
@@
As remarked above, by the $\Gamma$-structure, $G_{\varphi(e)}=A_{S_{0}}$ where $A_{S_{0}}\leq A_{\Gamma}$ is a parabolic subgroup. The partition of $K$ induces a partition of $V(\Gamma)=\bigcup_{i=0}^{n}S_{i}$, where for each $i$ and each $s\in S_{i}$ the subtree of $K$ fixed by $s$ is contained in $K_{i}\setminus\{\varphi(e)\}$. This union covers $V(\Gamma)$ because every generator is contained in a chunk (the corresponding parabolic fixes some vertex of $K$), and partitions because $\varphi(e)$ is fixed only by standard generators in $S_{0}$.
The tree structure gives rise to a decomposition of $A_{\Gamma}$ as an amalgamated product of factors $A_{S_{0}\cup S_{i}}$ (for $i\in\{1\dots n\}$) over the factor $A_{S_{0}}$. This is a visual splitting of the Artin group: if $s\in A_{S_{i}}$ and $t\in A_{S_{j}}$ are standard generators for $i\neq j$ both non-0, they act elliptically on $T$ and have disjoint fixed sets (separated by $\varphi(e)$, which neither fixes), so generate a free group by the ping-pong lemma.

### v2

@@ 3 Deformation spaces for Artin groups
In this section, given a large-type Artin group $A$, we construct a deformation space $\mathcal{D}$. Trees in this space will correspond to visual splittings of $A_{\Gamma}\cong A$ as amalgamated free products of chunks of $\Gamma$. These chunks lack separating simplices in the following sense. Note that in the following definition, $\mathrm{Star}$ is the open simplicial star.
@@
The second condition will help us pass to a finite index subgroup of $\mathrm{Out}(A)$, and find the point stabilisers in the deformation space we introduce. We expect this condition could be weakened slightly, but no weakening admits any new classes of Artin groups to be proven chunk rigid, and we expect this remains true in general.
@@
Note that since the graph of groups we built for $A$ is a tree, by Remark 2.13 the quotient of every $A$-tree in $\mathcal{D}_{\Gamma}$ is a tree, and the simplified theory of deformation spaces we considered in the previous section is sufficient for understanding this space of visual splittings.
@@
Consider the defining graph $\Gamma$ as in Figure 1, for simplicity with all edges labeled 7 (so that Proposition 3.6 applies). It has 4 chunks, coloured red, blue, yellow and green. The purple edge is shared by the red and blue chunks.
@@
If $A$ is connected, large-type Artin group with rigid chunks, with $A\cong A_{\Gamma}$. Then the construction of $\mathcal{D}_{\Gamma}$ did not depend on the choices in the construction of $T_{\Gamma}$, nor the choice of $\Gamma$ and isomorphism $A\cong A_{\Gamma}$.
@@
We repeatedly use Theorem 2.17.
Notice that $H\leq A_{\Gamma}$ fixes a point on $T_{\Gamma}$ if and only if it fixes a vertex, which occurs if and only if it is contained in some chunk parabolic subgroup of $A_{\Gamma}$. Immediately we see there is no dependence on the choices made in constructing $T_{\Gamma}$.
Suppose $A_{\Lambda}\cong A$ and fix an arbitrary isomorphism. We show that $T_{\Gamma}$ and $T_{\Lambda}$ are in the same deformation space. To see this simply note that $H$ (more precisely, its image in $A_{\Gamma}$) acts elliptically on $T_{\Gamma}$ if and only if it is contained in a chunk parabolic subgroup of $A_{\Gamma}$, which by Definition 3.2 occurs if and only if it is contained in a chunk parabolic subgroup of $A_{\Lambda}$, and this is true if and only if $H$ acts elliptically on $T_{\Lambda}$.
Finally, we wish to show $\mathrm{Out}(A)$ acts on $\mathcal{D}$. We define the action as follows: given $[\varphi]\in\mathrm{Out}(A)$ and $(T,\Omega)\in\mathcal{D}$, we declare $[\varphi]\cdot(T,\Omega)=(T,\Omega\circ\varphi)$.
We claim that $(T,\Omega\circ\varphi)\in\mathcal{D}$. Note $H$ fixes a point on $(T,\Omega)$ if and only $H$ is contained in a chunk parabolic subgroup of $A$ (this is well defined since $A$ is chunk rigid). This occurs if and only if $\varphi(H)$ is contained in a chunk parabolic subgroup of $A$, because images of chunk parabolic subgroups under automorphisms are chunk parabolic subgroups by chunk rigidity. Finally, we note that $\varphi(H)$ is contained in a chunk parabolic subgroup of $A$ if and only if $\varphi(H)$ acts elliptically on $(T,\Omega)$, if and only if $H$ acts elliptically on $(T,\Omega\circ\varphi)$.
Now, it is routine to check that the proposed action is a well defined action. ∎
We immediately apply Thoerem 2.23, and restrict our attention to the finite dimensional, $\mathrm{Out}(A)$ invariant subcomplex $\mathcal{F}\subseteq\mathcal{D}$ of surviving $A$-trees.
@@
It is enough to show that in the graph of groups we built for $A_{\Gamma}\curvearrowright T_{\Gamma}$, none of the inclusions of edge groups into vertex groups were equalities. Suppose $e=\{u,v\}$ was such that $G_{u}=G_{e}$, then this induces an inclusion $G_{u}\leq G_{v}$. However this induces an inclusion of the set of standard generators generating $G_{u}$ into those generating $G_{v}$. However the standard generators $G_{u}$ were supposed to span a subgraph which is a chunk, and this inclusion contradicts the maximality in the definition of a chunk.
@@
4 Type VF for outer automorphism groups of Artin groups
Our aim for this section is to show that, given $A$ an abstract connected large-type chunk rigid Artin group, $\mathrm{Out}(A)$ is type VF. From here, will will fix $A$ as an abstract connected large-type chunk rigid Artin group. We will write $\mathcal{F}$ for subcomplex of surviving trees in the canonical deformation space $\mathcal{D}$, as discussed in the previous section.
We will prove this in two parts. First we will show that $\mathcal{F}/\mathrm{Out}(A)$ is a finite complex. Secondly, we will show that $\mathrm{Out}(A)$ has a finite index subgroup that acts on $\mathcal{F}$ with type F simplex stabilisers.
@@ 4 Type VF for outer automorphism groups of Artin groups
Let $A$ be an abstract connected large-type chunk rigid Artin group. Then $\mathrm{Out}(A)$ is type VF.
@@
If $A$ is an even dihedral Artin group, then $\mathrm{Out}(A)$ is virtually $\mathbb{Z}$ [GHMR00], and the result follows. So assume $A$ is not an even dihedral Artin group, then by Lemma 4.14 the subgroup $\mathrm{Out}_{0,+}(A)$ defined in Section 4.2 is finite index in $\mathrm{Out}(A)$.
We consider the action of $\mathrm{Out}_{0,+}(A)$ on $\mathcal{F}$, which is contractible by Theorem 2.21 and Theorem 2.23. This action preserves the partial order on the vertices coming from elementary collapses, so is rigid.
By Proposition 4.11 there are finitely many $\mathrm{Out}(A)$ orbits of vertices in $\mathcal{F}$, and since $\mathrm{Out}_{0,+}(A)\leq_{f.i.}\mathrm{Out}(A)$ there are finitely many $\mathrm{Out}_{0,+}(A)$ orbits of vertices. Since the action is simplicial and $\mathcal{F}$ is finite dimensional, there are finitely many orbits of simplices, and $\mathcal{F}/\mathrm{Out}_{0,+}(A)$ is finite.
Moreover, by Lemma 4.18, every simplex stabiliser is of type F. It follows by Theorem 2.26 that $\mathrm{Out}_{0,+}(A)$ is of type F, and therefore $\mathrm{Out}(A)$ is of type VF. ∎
@@
Suppose $A_{\Gamma}$ is a connected large-type Artin group, and $\Gamma$ is either triangle-free, or has every label at least 6 (i.e. $A_{\Gamma}$ is XXXL-type). Then $\mathrm{Out}(A_{\Gamma})$ is of type VF.
@@ 4.1 Co-finiteness
It is convenient for us to regard $\Gamma$-trees as $A_{\Gamma}$-trees with distinguished fundamental domains, but one could equivalently see them as graph of groups decompositions of $A_{\Gamma}$ (over standard parabolic subgroups), with isomorphism being vertex and edge group preserving isomorphism of graphs of groups.
Lemma 4.5 justifies our definition of $\Gamma$-trees and isomorphism between them: if two trees in $\mathcal{F}$ can be given the structure of isomorphic $\Gamma$-trees (for the same $\Gamma$, so the definition of $\Gamma$-tree isomorphism makes sense), then they are in the same orbit under the action of $\mathrm{Out}(A)$.
We start with a technical lemma, which is essentially the observation that $\Gamma$-tree isomorphisms induce a graph-of-groups isomorphism on the $A_{\Gamma}$-trees, and hence there is an induced equivariant isometry.
@@
Suppose for $i\in\{1,2\}$ $(T_{i},\Omega_{i})$ are isomorphic $\Gamma$-trees (so $\Omega_{i}:A_{\Gamma}\rightarrow T_{i}$). Then they are equivariantly isometric: that is there is an isometry $\hat{f}:T_{1}\rightarrow T_{2}$ such that for all $g\in A_{\Gamma}$, $\hat{f}\circ\Omega_{1}(g)=\Omega_{2}(g)\circ\hat{f}.$
@@
For clarity, within the proof we will suppress reference to the actions $\Omega_{i}$ and interpret them implicitly. Our goal is to produce an isometry $\hat{f}:T_{1}\rightarrow T_{2}$ for all $g\in A_{\Gamma}$ and $x\in T_{1}$, $\hat{f}(gx)=g\hat{f}(x)$.
By the definition of $\Gamma$-tree isomorphism, there is an isometry between the fundamental domains of the $\Gamma$-trees, $f:K_{1}\rightarrow K_{2}$, which respects the edge and vertex groups. We define $\hat{f}:T_{1}\rightarrow T_{2}$ extending this isometry by such that for every point $gx\in T_{1}$ where $g\in A_{\Gamma}$ and $x\in K_{1}$ (all points can be written thus, since $K_{1}$ is a fundamental domain), $\hat{f}(gx):=gf(x)$. We need to check that this is a well defined, equivariant, and an isometry.
To see that $\hat{f}$ is well defined, note that if $gx=hy$ where $x,y\in K_{1}$, then it must be that $x=y$ since $K_{1}$ is a strict fundamental domain. Therefore $h^{-1}g\in G_{x}$. It follows that $h^{-1}g\in G_{f(x)}$, since $f$ is vertex group preserving, so $\hat{f}(gx)=gf(x)=hf(x)=\hat{f}(hx)$ as required.
It is clear that $\hat{f}$ sends any edge of $T_{1}$ isometrically to an edge of $T_{2}$. This is because any edge of $T_{1}$ can be written $ge$ where $g\in A_{\Gamma}$ and $e\subseteq K$, and $f$ is an isometry of $e$ onto an edge, so $\hat{f}$ is an isometry from the edge $ge$ to the edge $gf(e)$.
Therefore, to check that $\hat{f}$ is an isometric embedding, it is sufficient to check that it does not identify edges that share an endpoint. Suppose $e_{1}=\{u_{1},v\}$ and $e_{2}=\{v,u_{2}\}$ are distinct edges of $T_{1}$. If the edges are in distinct orbits, then it is not hard to check, using that fact that $K_{1}$ is a strict fundamental domain, that there is $g\in A_{\Gamma}$ such that $g^{-1}(e_{1}\cup e_{2})\subseteq K_{1}$, and the required result follows from the fact that $f$ is an isometric embedding. Otherwise, suppose $h^{-1}e_{1}\subseteq K_{1}$, and so $h^{-1}u_{2}=gh^{-1}u_{1}$ where $g\in G_{h^{-1}v}\setminus G_{h^{-1}u_{1}}$ by assumption that $u_{1}$ and $u_{2}$ are distinct. Note that
$G_{h^{-1}v}\setminus G_{h^{-1}u_{1}}=G_{f(h^{-1}v)}\setminus G_{f(h^{-1}u_{1})}=G_{h^{-1}\hat{f}(v)}\setminus G_{h^{-1}\hat{f}(u_{1})}=h^{-1}(G_{\hat{f}(v)}\setminus G_{\hat{f}(u_{1})}),$
where the crucial equality is the first where we use that $f$ respects the vertex groups. In particular,
$hgh^{-1}\in G_{\hat{f}(v)}\setminus G_{\hat{f}(u_{1})}$
@@
$\hat{f}(u_{2})=\hat{f}(hgh^{-1}u_{1})=hgf(h^{-1}u_{1})=hgh^{-1}\hat{f}(u_{1}),$
and $hgh^{-1}\in G_{\hat{f}(v)}$, so $hgh^{-1}$ takes $\hat{f}(e_{1})$ to $\hat{f}(e_{2})$, but $hgh^{-1}\notin G_{\hat{f}(u_{2})}$, so these edges are distinct as required.
That $\hat{f}$ is surjective follows immediately from the fact any point in $T_{2}$ is in a translate of the fundamental domain, so can be written as $gf(x)=\hat{f}(gx)$.
Finally, equivariance of $\hat{f}$ follows almost immediately from the definition. ∎
@@
For $i\in\{1,2\}$ let $(T_{i},\Omega_{i})$ be two $A$-trees (so $\Omega_{i}:A\rightarrow\mathrm{Aut}(T_{i})$) in $\mathcal{F}$. Suppose $\Gamma$ is an abstract defining graph, and there are isomorphisms $\psi_{i}:A_{\Gamma}\rightarrow A$. Suppose that the $A_{\Gamma}$-trees $(T_{i},\Omega_{i}\circ\psi_{i})$ can be seen as isomorphic $\Gamma$-trees (by choice of fundamental domains $K_{i}$).
@@
We claim that $K=\mathrm{Hull}(\{v_{i}\}_{i\in I})$. To see this, suppose for contradiction there was a nonempty subtree of $K$ outside of this hull, attached to a vertex $v\in\mathrm{Hull}(\{v_{i}\}_{i\in I})$ by an edge $e$. Let $G_{v}=A_{S}$ (we make no assumptions beyond it being standard parabolic, which follows from the definition of a $\Gamma$-tree). Now take $w$ an arbitrary edge or vertex of this extremal subtree. By definition of a $\Gamma$-tree, $w$ has stabiliser $A_{S^{\prime}}$ where $S^{\prime}$ is either a standard generator, or a pair of standard generators spanning an spherical dihedral parabolic subgroup. Take $s\in S^{\prime}$. This generator, viewed as a vertex in $\Gamma$, occurs in a chunk of $\Gamma$. Therefore it fixes the unique vertex $v_{i}$ in $K$ corresponding to this chunk, and the geodesic from $w$ to $v_{i}$. Given that $w$ is a vertex in an extremal subtree, this geodesic passes through $v$. It follows that $s\in A_{S}$. Since $s$ and $w$ were generic, it follows that the stabiliser of every point in the extremal subtree has stabiliser including into $A_{S}$. This contradicts minimality of the action, completing the proof of the claim.
Now we show that there is a uniform bound on the number of vertices in $K$ depending only on $\Gamma$. Since $K=\mathrm{Hull}(\{v_{i}\}_{i\in I})$, we can write $K$ as a union of $|I|^{2}$ geodesics, in particular $K=\bigcup_{i,j\in I}[v_{i},v_{j}]$. Note that $|I|$ is the number of chunks, and depends only on $\Gamma$.
@@
Now suppose $v$ is internal to $\mu$ with valence exactly 2 in $K$, and the call $e_{1},e_{2}$ the adjacent edges. One case is that $v\in\{v_{k}\}_{k\in I\setminus\{i,j\}}$, but this happens at most $|I|-2$ times. Thus we focus on the other case, so $G_{v}=A_{S}$, where $S$ is either a single standard generator, or two standard generators spanning an spherical dihedral parabolic subgroup.
@@
For expansion, we may have to change $\Gamma$ to some new $\Gamma^{\prime}$ (with $A_{\Gamma^{\prime}}\cong A_{\Gamma}$). This case is more delicate than the previous one, with the tricky part being picking the right $\Gamma^{\prime}$ and fundamental domain on $T^{\prime}$ (once we have done so, checking there is a $\Gamma^{\prime}$-tree structure is laborious but not difficult).
@@
Suppose that $T,T^{\prime}\in\mathcal{F}$ are both surviving, and $\varphi:T^{\prime}\rightarrow T$ is a collapse of an edge $e=\{u,v\}$ with $G_{u}\leq G_{v}$. Suppose that $T$ is a $\Gamma$-tree. Then there is $\Gamma^{\prime}$ and an isomorphism $\psi:A_{\Gamma^{\prime}}\rightarrow A_{\Gamma}$, such that $T^{\prime}$, when viewed as an $A_{\Gamma^{\prime}}$-tree (by precomposing the action $A_{\Gamma}\rightarrow\mathrm{Aut}(T^{\prime})$ with $\psi$) is a $\Gamma^{\prime}$-tree.
@@
As remarked above, by the $\Gamma$-structure, $G_{\varphi(e)}=A_{S_{0}}$ where $A_{S_{0}}\leq A_{\Gamma}$ is a parabolic subgroup. The partition of $K$ induces a partition of $V(\Gamma)=\bigcup_{i=0}^{n}S_{i}$, where for each $i\neq 0$ and each $s\in S_{i}$ the subtree of $K$ fixed by $s$ is contained in $K_{i}\setminus\{\varphi(e)\}$. This union covers $V(\Gamma)$ because every generator is contained in a chunk (the corresponding parabolic fixes some vertex of $K$), and partitions because $\varphi(e)$ is fixed only by standard generators in $S_{0}$.
The tree structure gives rise to a decomposition of $A_{\Gamma}$ as an amalgamated product of factors $A_{S_{0}\cup S_{i}}$ (for $i\in\{1\dots n\}$) over the factor $A_{S_{0}}$. This is a visual splitting of the Artin group. To see this, take $i,j\in\{1\dots n\}$ distinct, and $s\in A_{S_{i}}$ and $t\in A_{S_{j}}$ standard generators. Both $s$ and $t$ act elliptically on $T$ and have disjoint fixed sets (separated by $\varphi(e)$, which neither $s^{k}$ nor $t^{k}$ fixes for any $k\neq 0$), so generate a free group by the ping-pong lemma.